An initially stationary 2.0kg object accelerates horizontally and uniformly to a speed of 10m/s in 3.0s. In that 3.0s interval, how much work is done on the object by the force accelerating it? What is the instantaneous power due to that force at the end of the interval and at the end of the first half of the interval?
Answer #1
a = Δv / Δt
a = 10 / 3
a = 3.333333 m/s2
F = ma
F = 3.333(2)
F = 6.666 N
W = Fd
W = 6.666(.5at2)
W = 6.666(.5(3.333)(9))
W = 100 J
The instantaneous power at ANY instant is Force *Velocity:
At the end: 6.6666(10) = 67 Watts.

Halfway: at 1.5 secs: Power = 33 Watts